OpenStudy (anonymous):

B is between A and C. AB=2x-4 BC=7x+5 AC=82 What is the value of x? Show your work for credit.

OpenStudy (amistre64):

does AB + BC = AC?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then all thats left is to fill in the peices and solve for x

OpenStudy (amistre64):

AB + BC = AC (2x-4) + (7x+5 ) = 82

OpenStudy (anonymous):

how do i find x tho?

OpenStudy (amistre64):

by combing like terms to begin with

OpenStudy (amistre64):

since there is a + in the middle, the paranthesis are really useless 2x-4 + 7x+5 = 82